CH/Agen/93/03

Agenda

Charissa Group Meeting

To be held on: Thursday, 13 May 1993, 11.15 am To be held at: Birmingham

Distribution: WNC,NMC,RAC,BRF,GJG,SJAH,RAH,JK,JSL,ACM, VP,WDMR,GT,RPW,DDW,DLW,JZ , all by e-mail/server

1. Minutes of Meeting of 7 April 1993
2. Matters Arising/ Actions
3. Update on NSC/SERC matters  (DLW)
4. Progress on detector tests  (DLW)
5. Progress on chamber & detector assemblies  (JSL)
6. Progress on gas supply  (RAC)
7. Progress on electronics & data acquisition  (RAH)
8. Developments on event manager & GEC emulator  (VP)
9. Requirements for hybrid detector operation at the ANU (RAC)
10. Discussion on trip to Strasbourg
11. Agree final timetable for ANU
12. Status of shipment to the ANU  (DLW)
13. Documentation for ANU Running
14. Status of papers and need for publications
15. Any other business
16. Date of next meeting (15 July 1993 proposed)

W.N. Catford (Secretary) 4 May 1993